# Representation Retrieval Learning for Heterogeneous Data Integration

### v1

@@ 1.1 Related works
Distribution heterogeneity has attracted much attention in data integration. Primarily, it is important to discover the information sharing structure which enables one to capture common information across data sources. In particular, the joint-and-individual structure [26, 45] has been widely adopted, assuming a jointly shared component by all data sources and individual components unique to each data source. This structure is also applied to latent factor regression [51], canonical correlation analysis [25] and multi-task linear regression [9]. Recently, the partially sharing structure was developed [14, 27, 58] to accommodate more flexible information sharing structure. Specifically, each component can be shared by any number of data sources, which empower a greater heterogeneity level. Indeed, this information sharing structure is more common and natural in real applications (see motivating examples in Appendix 1.2). There are two techniques to discover a partially sharing structure: first, partially sharing structure can be encoded by binary matrices, which can be estimated via forward selection [27], or two-step selection [14]. Both estimations are computationally infeasible for high-dimensional or large-scale data. Second, a sparse-induced penalty [48, 40] is adopted to recover the partially sharing structure [30, 23, 58], which reduces the estimation complexity due to its convexity.
@@ 1.2 Motivating example
In order to examine the homogeneity of the marginal distribution of each modality, we apply sparse PCA (Witten et al., [56], sparsity is necessary here because the dimension of covariates is greater than sample sizes for all modalities and data sources) to project each modality into 2-dimensional space; and the density of each data source can be found in Appendix LABEL:app_sec:_motivation_figures. The density plots of the MRI modality have apparent discrepancy across the first three data sources, which suggests potential covariate shift.
Next, we examine the posterior homogeneity across four data sources. Since the true models are unknown to us and [63, 60] adopted a linear model for their analysis, we calculate the marginal correlation between each covariate and response across four sources, and evaluate whether these marginal correlations are homogeneous across sources. Figure LABEL:fig:_adni_modalities_marginal_correlation in Appendix LABEL:app_sec:_motivation_figures presents the plots of marginal correlations for covariates among three modalities, respectively. In the MRI and Gene modalities, marginal correlations across data sources are quite heterogeneous, where many features exhibit opposite marginal correlation to responses across data sources. Therefore, posterior heterogeneity should be considered when analyzing this dataset.
@@ 3.1 Representation Retrieval Learning
We consider the representation dictionary as a set of representers, denoted as $\Theta=\{\theta_{1},\cdots,\theta_{D}\}$. Formally, we define a representer as a univariate function mapping $\theta_{d}:\mathcal{X}\rightarrow\mathbb{R}$, so we use $\theta_{d}$ and $\theta_{d}(\mathbf{x})$ interchangeably in later discussion. Multivariate representer outputs are also allowed, where the related extension can be found in Appendix LABEL:app_sec:_multi-dim-rep. For each task, we assume that a subset of informative representers from $\Theta$ are retrieved, denoted as $\Gamma^{(s)}=\{\theta_{d_{1}},\cdots,\theta_{d_{s}}\}$. Naturally, we define $\Gamma^{(s)}(\mathbf{x})=(\theta_{d_{1}}(\mathbf{x}),\cdots,\theta_{d_{s}}(\mathbf{x})):\mathcal{X}\rightarrow\mathbb{R}^{|\Gamma^{(s)}|}$. Since we do not impose any structural assumptions on $\Gamma^{(s)}$, both joint-and-individual or partially sharing structures are incoporated as shown in Figure 3.
@@ 4.1 Excess risk bound for $R^{2}$ learning
The quantity $\mathbb{E}_{\boldsymbol{\sigma}}\mathbf{G}_{\beta}(\boldsymbol{\sigma})$, analogous to the Rademacher complexity, quantifies the capacity of the representation dictionary to fit random noise for any fixed $\beta\in\mathcal{H}$. A similar measure has been studied for linear dictionaries in [30]. In Lemma 4.1, we refine the analysis by leveraging the support of the given $\beta^{(s)}$. See Appendix LABEL:app_sec:_proof for technical details. The upper bound (4.1) only considers the $\ell_{1}$ penalty on $\beta^{(s)}$’s where $\breve{\gamma}(\tau)$ and $\bar{g}(\tau)$ are functions of $\beta$, which provides a precise description of the role of “integrativeness” of representers in average generalization error bound. In contrast, the upper bound (4.2) is an uniform bound for any $\beta\in\mathcal{H}$, which is useful to establish the results in the following Theorem 4.1.
@@ 5.1 Simulations for Multi-task Data
We also compare our method against prominent multi-task methods from the literature. First, we consider Multi-task Representation Learning (MTRL; Du et al., 8, Tripuraneni et al., 50), which employs a uniform representer along with source-specific learners to accommodate both linear and nonlinear patterns. Second, we evaluate a penalized ERM algorithm (pERM; Tian et al., 47), which allows for similar, though not identical, linear representations across sources and accommodates a subset of outlier sources. Third, Adaptive and Robust Multi-task Learning (ARMUL; Duan and Wang, 9) is included, as it encourages regression coefficients to conform to a low-rank representation, except for a small set of outliers. pERM is implemented using the open-source code provided by [47] (https://github.com/ytstat/RL-MTL-TL), and ARMUL is implemented based on the code from [9] (https://github.com/kw2934/ARMUL?tab=readme-ov-file). IL (both linear and neural network versions) and MTRL are implemented in pytorch. We adopt default algorithm setups, such as learning rate and epochs, in all the above implementation, and tune the penalty coefficients and neural nets structures in all methods with optuna.
@@
where covariates $\mathbf{X}^{(s)}$ follow a multivariate standard normal distribution with dimension $p=30$. The representer dictionary $\Theta(\cdot)=(\theta_{1},\cdots,\theta_{30})$ contains 30 representers, including 6 linear representers and 24 nonlinear representers, and detailed specification of these representers can be found in Appendix LABEL:app_sec:_spec_R2_sim. For each source, only three representers are informative, i.e., $\lVert\beta^{(s)}\rVert_{0}=3$ for all $s=1,\cdots,S$. The noise term $\epsilon^{(s)}$ are i.i.d samples of standard normal distribution.
@@ 5.2 Simulations for Multi-modality Block-mising Data
where modality-specific covariates $\mathbf{X}_{m}^{(s)}$ are sampled from a $q=40$-dimensional multivariate normal distribution with modality-specific covariance structures (details are provided in Appendix LABEL:app_sec:_spec_BR2_sim). For fair comparison, we restrict the representers to the linear case (orthogonal bases in $\mathbb{R}^{q}$) as DISCOM is designed for linear models. For each modality, only two representers are active, i.e., $\lVert\beta_{m}^{(s)}\rVert_{0}=2$, and the noise term $\epsilon^{(s)}$ are i.i.d samples of standard normal distribution..

### v2

@@ 1.1 Related works
Distribution heterogeneity has attracted much attention in data integration. Primarily, it is important to discover the information sharing structure which enables one to capture common information across data sources. In particular, the joint-and-individual structure [27, 46] has been widely adopted, assuming a jointly shared component by all data sources and individual components unique to each data source. This structure is also applied to latent factor regression [53], canonical correlation analysis [26] and multi-task linear regression [9]. Recently, the partially sharing structure was developed [14, 28, 60] to accommodate more flexible information sharing structure. Specifically, each component can be shared by any number of data sources, which empower a greater heterogeneity level. Indeed, this information sharing structure is more common and natural in real applications (see motivating examples in Section 1.2). There are two techniques to discover a partially sharing structure: first, partially sharing structure can be encoded by binary matrices, which can be estimated via forward selection [28], or two-step selection [14]. Both estimations are computationally infeasible for high-dimensional or large-scale data. Second, a sparse-induced penalty [49, 41] is adopted to recover the partially sharing structure [31, 24, 60], which reduces the estimation complexity due to its convexity.
@@ 1.2 Motivating example
In order to examine the homogeneity of the marginal distribution of each modality, we apply sparse PCA (Witten et al., [58], sparsity is necessary here because the dimension of covariates is greater than sample sizes for all modalities and data sources) to project each modality into 2-dimensional space; and the density of each data source can be found in Appendix C. The density plots of the MRI modality have apparent discrepancy across the first three data sources, which suggests potential covariate shift.
Next, we examine the posterior homogeneity across four data sources. Since the true models are unknown to us and [65, 62] adopted a linear model for their analysis, we calculate the marginal correlation between each covariate and response across four sources, and evaluate whether these marginal correlations are homogeneous across sources. Figure 12 in Appendix C presents the plots of marginal correlations for covariates among three modalities, respectively. In the MRI and Gene modalities, marginal correlations across data sources are quite heterogeneous, where many features exhibit opposite marginal correlation to responses across data sources. Therefore, posterior heterogeneity should be considered when analyzing this dataset.
@@ 3.1 Representation Retrieval Learning
We consider the representation dictionary as a set of representers, denoted as $\Theta=\{\theta_{1},\cdots,\theta_{D}\}$. Formally, we define a representer as a univariate function mapping $\theta_{d}:\mathcal{X}\rightarrow\mathbb{R}$, so we use $\theta_{d}$ and $\theta_{d}(\mathbf{x})$ interchangeably in later discussion. Multivariate representer outputs are also allowed, where the related extension can be found in Appendix A. For each task, we assume that a subset of informative representers from $\Theta$ are retrieved, denoted as $\Gamma^{(s)}=\{\theta_{d_{1}},\cdots,\theta_{d_{s}}\}$. Naturally, we define $\Gamma^{(s)}(\mathbf{x})=(\theta_{d_{1}}(\mathbf{x}),\cdots,\theta_{d_{s}}(\mathbf{x})):\mathcal{X}\rightarrow\mathbb{R}^{|\Gamma^{(s)}|}$. Since we do not impose any structural assumptions on $\Gamma^{(s)}$, both joint-and-individual or partially sharing structures are incoporated as shown in Figure 3.
@@ 4.1 Excess risk bound for $R^{2}$ learning
The quantity $\mathbb{E}_{\boldsymbol{\sigma}}\mathbf{G}_{\beta}(\boldsymbol{\sigma})$, analogous to the Rademacher complexity, quantifies the capacity of the representation dictionary to fit random noise for any fixed $\beta\in\mathcal{H}$. A similar measure has been studied for linear dictionaries in [31]. In Lemma 4.1, we refine the analysis by leveraging the support of the given $\beta^{(s)}$. See Appendix D for technical details. The upper bound (4.1) only considers the $\ell_{1}$ penalty on $\beta^{(s)}$’s where $\breve{\gamma}(\tau)$ and $\bar{g}(\tau)$ are functions of $\beta$, which provides a precise description of the role of “integrativeness” of representers in average generalization error bound. In contrast, the upper bound (4.2) is an uniform bound for any $\beta\in\mathcal{H}$, which is useful to establish the results in the following Theorem 4.1.
@@ 5.1 Simulations for Multi-task Data
We also compare our method against prominent multi-task methods from the literature. First, we consider Multi-task Representation Learning (MTRL; Du et al., 8, Tripuraneni et al., 51), which employs a uniform representer along with source-specific learners to accommodate both linear and nonlinear patterns. Second, we evaluate a penalized ERM algorithm (pERM; Tian et al., 48), which allows for similar, though not identical, linear representations across sources and accommodates a subset of outlier sources. Third, Adaptive and Robust Multi-task Learning (ARMUL; Duan and Wang, 9) is included, as it encourages regression coefficients to conform to a low-rank representation, except for a small set of outliers. pERM is implemented using the open-source code provided by [48] (https://github.com/ytstat/RL-MTL-TL), and ARMUL is implemented based on the code from [9] (https://github.com/kw2934/ARMUL?tab=readme-ov-file). IL (both linear and neural network versions) and MTRL are implemented in pytorch. We adopt default algorithm setups, such as learning rate and epochs, in all the above implementation, and tune the penalty coefficients and neural nets structures in all methods with optuna.
@@
where covariates $\mathbf{X}^{(s)}$ follow a multivariate standard normal distribution with dimension $p=30$. The representer dictionary $\Theta(\cdot)=(\theta_{1},\cdots,\theta_{30})$ contains 30 representers, including 6 linear representers and 24 nonlinear representers, and detailed specification of these representers can be found in Appendix B.1. For each source, only three representers are informative, i.e., $\lVert\beta^{(s)}\rVert_{0}=3$ for all $s=1,\cdots,S$. The noise term $\epsilon^{(s)}$ are i.i.d samples of standard normal distribution.
@@ 5.2 Simulations for Multi-modality Block-mising Data
where modality-specific covariates $\mathbf{X}_{m}^{(s)}$ are sampled from a $q=40$-dimensional multivariate normal distribution with modality-specific covariance structures (details are provided in Appendix B.2). For fair comparison, we restrict the representers to the linear case (orthogonal bases in $\mathbb{R}^{q}$) as DISCOM is designed for linear models. For each modality, only two representers are active, i.e., $\lVert\beta_{m}^{(s)}\rVert_{0}=2$, and the noise term $\epsilon^{(s)}$ are i.i.d samples of standard normal distribution..